OpenStudy (anonymous):

HELP PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1 help solving systems of equations with three variables 1. 2a+c=-10, b-c=15, a-2b+c=-5 2. x+y+z=3, 13x+2z=2, -x-5z=-5

OpenStudy (anonymous):

x+y+z=3, 13x+2z=2, -x-5z=-5 \[x+y+z=3\] \[13x+0y+2z=2\] \[-x+0y+-5z=-5\] you can use various methods to solve this system of equation. I'm assuming you have learned either substitution or elimination. To eliminate, you can add the second equation with the third equation multiplied by 13. \[13x+0y+2z=2\] \[(13*(-x+0y+-5z=-5)=>-13x+0y-65z=-65\] If you add the columes, you'll find that you'll cancel the x values. You should get something like \[0x+0y-63z=-63\] If you solve for this equation, you'll find that z=1 With that solved, you can substitute the z=1 back into either the second or third equation to find x. Then do the same to find y in the first equation.

OpenStudy (anonymous):

ok thx